# Search for QCD Axions in light of String Theory

### v1

@@ Abstract
The QCD axion stands as one of the most promising candidates for resolving the strong CP problem. However, the value of the axion’s decay constant $f_{a}$ and, by extension, its mass $m_{a}$, remain uncertain within the framework of effective field theory, posing a challenge for experimental detection. Fortunately, fields such as cosmology and astrophysics can offer crucial clues about potential mass ranges. Additionally, string theory and the more recent swampland principles might shed some light on the subject. The most straightforward string theory compactifications suggest that $f_{a}$ is around the GUT scale, leading to an anthropic abundance of dark matter. We found that this range can be detected through hydrogen atomic transitions. The recent concept of the dark dimension scenario introduces an alternative possibility. If axions are confined to the four-dimensional Standard Model brane, their decay constant $f_{a}$ would be on the order of $10^{10}$ GeV. In this scenario, where axions constitute only a minor portion of dark matter, we show that a laser-interferometry setup would be an effective detection method.
@@ I Introduction
The uncertainty of the QCD axion primarily stems from the axion decay constant $f_{a}$. The earliest model Weinberg (1978) assumed it to be at the electro-weak scale or approximately 250 GeV, but experiments have ruled out this scenario. Subsequent models Kim (1979); Shifman et al. (1980); Zhitnitsky (1980); Dine et al. (1981) separate the electro-weak scale from the Peccei-Quinn scale, assuming a much higher scale. From the perspective of effective field theory, the uncertainty is quite large. Current laboratory searches and astrophysical constraints typically exclude values of $f_{a}$ below $10^{8}$ GeV Sikivie (1983); Anastassopoulos et al. (2017).
@@
String theory and the recent dark dimension scenario offer insights into the axion/axion-like particle decay constant $f_{a}$vrcek and Witten (2006); Montero et al. (2023); Gendler and Vafa (2024). Combined with the potential of QCD instantons, the QCD axion mass can be fixed within a relatively narrow range. The remainder of this paper is organized as follows: In Section II, we demonstrate that the hydrogen 1S triplet state is suitable for detecting QCD axions in the general string compactification scenario. In Section III, we show that the interferometry scheme is appropriate for detecting QCD axions in the dark dimension scenario.
@@ II.1 Theoretical Considerations for $f_{a}$ in String Compactifications
When the value of $f_{a}$ is much greater than $10^{12}$ GeV, one must consider whether the resulting anthropic dark matter window is consistent with the isocurvature perturbations of the cosmic microwave background (CMB) radiation Hertzberg et al. (2008). In this case, the dark matter axion field has an initial field value $a=\theta f_{a}$ before cosmological inflation. The Gibbons-Hawking temperature during inflation leads to a small standard deviation $\sigma_{\theta}f_{a}$ in the axion field, given by
@@ III.2 Detecting Dark Dimension Scenario Axions by Interferometry
The utilization of interferometers, however, comes with the presence of shot noise, which reduces the sensitivity to detecting axions. For a laser beam comprising $N$ incoming photons, the shot noise is $\sqrt{N}$. Consequently, the signal-to-noise ratio is reduced to $(g_{a\gamma\gamma}BL)^{2}N/\sqrt{N}$. This converts to a sensitivity of axion coupling $g_{a\gamma\gamma}<(BL)^{-1}N^{-1/4}$. An implementation of squeezed light can assist in mitigating shot noises. When the interferometer operates near a dark fringe, the shot noise can be theoretically reduced indefinitely. Implementations of squeezed light are demonstrated with a $15$ dB shot noise reduction Vahlbruch et al. (2016). Additionally, by adding optical delay lines or Fabry-Perot (FP) cavities, the signal can be further enhanced by a factor of $n$, where $n$ is the number of times the laser beam is repeated inside the device.
By employing $n\sim 10^{4}$, $B\sim 10$ T, $L\sim 10$ m with a $10$ W ($\lambda=1\mu$ m) laser, after 10 days of operation, the experiment can probe axions with $g_{a\gamma\gamma}\geq 8.4\times 10^{-11}$ GeV-1 to a $5\sigma$ significance level, a purely laboratory experiment probing axion-photon coupling at a level competitive to limits from solar axion searches. If one also uses a squeezed-light laser of order ${\cal O}(10)$ dB shot noise reduction, the exclusion limit could reach $f_{a}\sim 10^{9}-10^{10}$ GeV, which is the preferred region in the dark dimension scenario.

### v2

@@ Abstract
The QCD axion stands as one of the most promising candidates for resolving the strong CP problem. However, the value of the axion’s decay constant $f_{a}$ and, by extension, its mass $m_{a}$, remain uncertain within the framework of effective field theory, posing a challenge for experimental detection. Fortunately, fields such as cosmology and astrophysics can offer crucial clues about potential mass ranges. Additionally, string theory and the more recent swampland principles might shed some light on the subject. The most straightforward string theory compactifications suggest that $f_{a}$ is around the GUT scale, leading to a quantum abundance of dark matter. We found that this range can be detected through hydrogen atomic transitions. The recent concept of the dark dimension scenario introduces an alternative possibility. If axions are confined to the four-dimensional Standard Model brane, their decay constant $f_{a}$ would be on the order of $10^{10}$ GeV. In this scenario, where axions constitute only a minor portion of dark matter, we show that a laser-interferometry setup would be an effective detection method.
@@ I Introduction
The uncertainty of the QCD axion primarily stems from the axion decay constant $f_{a}$. The earliest model Weinberg (1978) assumed it to be at the electro-weak scale or approximately 250 GeV, but experiments have ruled out this scenario. Subsequent models Kim (1979); Shifman et al. (1980); Zhitnitsky (1980); Dine et al. (1981) separate the electro-weak scale from the Peccei-Quinn scale, assuming a much higher scale. From the perspective of effective field theory, the uncertainty is quite large. Current laboratory searches and astrophysical constraints typically exclude values of $f_{a}$ below $10^{8}$ GeV Sikivie (1983); Anastassopoulos et al. (2017). For a more recent review, see Ref.Ringwald (2024).
@@
String theory and the recent dark dimension scenario offer insights into the axion/axion-like particle decay constant $f_{a}$vrcek and Witten (2006); Visinelli and Vagnozzi (2019); Montero et al. (2023); Gendler and Vafa (2024). Combined with the potential of QCD instantons, the QCD axion mass can be fixed within a relatively narrow range. The remainder of this paper is organized as follows: In Section II, we demonstrate that the hydrogen 1S triplet state is suitable for detecting QCD axions in the general string compactification scenario. In Section III, we show that the interferometry scheme is appropriate for detecting QCD axions in the dark dimension scenario.
@@ II.1 Theoretical Considerations for $f_{a}$ in String Compactifications
When the value of $f_{a}$ is much greater than $10^{12}$ GeV, one must consider whether the resulting dark matter window is consistent with the isocurvature perturbations of the cosmic microwave background (CMB) radiation Hertzberg et al. (2008); Visinelli and Gondolo (2009). In this case, the dark matter axion field has an initial field value $a=\theta f_{a}$ before cosmological inflation. The Gibbons-Hawking temperature during inflation leads to a small standard deviation $\sigma_{\theta}f_{a}$ in the axion field, given by
@@ II.2 Detecting GUT Scale Dark Matter Axions with Hydrogen Atoms
Figure 3: The mass range that can be covered is determined by the available Zeeman field strength. To cover Region I ($m_{a}\in[10^{-5}\,\text{eV},\,2\times 10^{-5}\,\text{eV}]$), a field strength of 0.15 Tesla is needed. Assuming 10% of this region is covered in a one-year data-taking period using 1 mole of H atoms, the sensitivity is $f_{a}=5.8\times 10^{11}\,\text{GeV}$. For Region II ($m_{a}\in[10^{-8}\,\text{eV},\,2\times 10^{-8}\,\text{eV}]$), a field strength of 0.002 Tesla is required. Assuming 10% of this region is covered using $10^{3}$ moles of H over a one-year data-taking period, the sensitivity is $f_{a}=5.8\times 10^{14}\,\text{GeV}$. The yellow band defines the QCD axion parameter space. The solid-colored regions, SN1987A and Solar Core, have been excluded by experiments and observations. The light coral region can be probed using this hydrogen splitting method if a 1 Tesla Zeeman field and an adequate amount of hydrogen atoms are deployed.
@@ III.2 Detecting Dark Dimension Scenario Axions by Interferometry
Figure 5: Assuming $n\sim 10^{6}$, $B\sim 40$ T, $L\sim 30$ m with a 10 W ($\lambda=1\,\mu$m) laser, and squeezed light achieving approximately ${\cal O}(10)$ dB shot noise reduction, the laser experiment can, after 10 days of operation, probe axions with $g_{a\gamma\gamma}\geq 2.3\times 10^{-13}\,\text{GeV}^{-1}$ at a $5\sigma$ significance level.
@@
The utilization of interferometers, however, comes with the presence of shot noise, which reduces the sensitivity to detecting axions. For a laser beam comprising $N$ incoming photons, the shot noise is $\sqrt{N}$. Consequently, the signal-to-noise ratio is reduced to $(g_{a\gamma\gamma}BL)^{2}N/\sqrt{N}$. This converts to a sensitivity of axion coupling $g_{a\gamma\gamma}<(BL)^{-1}N^{-1/4}$. An implementation of squeezed light can assist in mitigating shot noises. When the interferometer operates near a dark fringe, the shot noise can be theoretically reduced indefinitely. Implementations of squeezed light are demonstrated with a $15$ dB shot noise reduction Vahlbruch et al. (2016). Additionally, by adding optical delay lines or Fabry-Perot (FP) cavities, the signal can be further enhanced by a factor of $n$, where $n$ is the number of times the laser beam is repeated inside the device. Equivalently, this results in an $n^{0.5}$ improvement in our ability to constrain $g_{a\gamma\gamma}$.
By employing $n\sim 10^{6}$, $B\sim 40$ T, $L\sim 30$ m with a 10 W ($\lambda=1\,\mu$m) laser, and squeezed light achieving approximately ${\cal O}(10)$ dB shot noise reduction, the experiment can, after 10 days of operation, probe axions with $g_{a\gamma\gamma}\geq 2.3\times 10^{-13}\,\text{GeV}^{-1}$ at a $5\sigma$ significance level.